Find the amount of interest that a person will earn if he invests an amount of R9000 for 30 months, at an interest rate of 6% per annum, with the interest compounded semi-annually. (Leave your answer as an expression in exponential form as you may not use a calculator.)
1
Expert's answer
2019-08-15T10:05:40-0400

Compound interest="P(1+\frac{r}{n})^{nt}-P"

="9000\times(1+\frac{0.06}{2})^{2\times{2.5}}-9000"